# Trans-Glasso: A Transfer Learning Approach to Precision Matrix Estimation

### v1

@@ Appendix B Preliminary Lemmas
Definition B.1 (Sub-Gaussian Random Variable, Definition 2 of Ravikumar et al. [2011]).
@@
Assume that we obtain samples from $K$ distributions, each with a mean of zero and a covariance matrix of $\Sigma^{(k)}$. Let $\{X^{(k)}_{i}=(X^{(k)}_{i1},\ldots,X^{(k)}_{id})^{\top}\}_{i=1}^{n_{k}}$ represent $n_{k}$ independently distributed samples from the $k$-th distribution. In addition, we assume that $X^{(k)}_{ij}/\sqrt{\Sigma^{(k)}_{jj}}$ is sub-Gaussian with parameter $\sigma$ as defined in Definition B.1. Let
@@
Let $\widebar{X}^{(k)}_{ij}={X^{(k)}_{ij}}/{\sqrt{\Sigma^{(k)}_{jj}}}$, $i=1,\ldots,n_{k}$, $k=1,\ldots,K$, $j=1,\ldots,d$. Define $U^{(k)}_{i,jl}=\widebar{X}^{(k)}_{ij}+\widebar{X}^{(k)}_{il}$, $V^{(k)}_{i,jl}=\widebar{X}^{(k)}_{ij}-\widebar{X}^{(k)}_{il}$, and $\rho^{(k)}_{jl}={\Sigma^{(k)}_{jl}}/{\sqrt{\Sigma^{(k)}_{jj}\cdot\Sigma^{(k)}_{ll}}}$. Then it follows from Proposition 2.9 and (2.18) in Wainwright [2019], that
@@
$\mathbb{E}$ $\left[\exp\left\{\lambda\left[\sum^{n_{k}}_{i=1}\left(\widebar{X}^{(k)}_{ij}\widebar{X}^{(k)}_{il}-\rho^{(k)}_{jl}\right)\right]\right\}\right]$
$=\mathbb{E}\left[\exp\left\{\frac{\lambda}{4}\left[\sum^{n_{k}}_{i=1}\left(\left(U^{(k)}_{i,jl}\right)^{2}-2\left(1+\rho^{(k)}_{jl}\right)\right)\right]\right\}\cdot\exp\left\{-\frac{\lambda}{4}\left[\sum^{n_{k}}_{i=1}\left(\left(V^{(k)}_{i,jl}\right)^{2}-2\left(1-\rho^{(k)}_{jl}\right)\right)\right]\right\}\right]$
@@
$\mathbb{E}$ $\left[\exp\left\{\lambda\left[\sum^{n_{k}}_{i=1}\left(X^{(k)}_{ij}X^{(k)}_{il}-\Sigma^{(k)}_{jl}\right)\right]\right\}\right]$
$=\mathbb{E}\left[\exp\left\{\lambda\sqrt{\Sigma^{(k)}_{jj}}\sqrt{\Sigma^{(k)}_{ll}}\left[\sum^{n_{k}}_{i=1}\left(\widebar{X}^{(k)}_{ij}\widebar{X}^{(k)}_{il}-\rho^{(k)}_{jl}\right)\right]\right\}\right]$
$\leq\mathbb{E}\left[\exp\left\{\lambda|\Sigma^{(k)}|_{\infty}\left[\sum^{n_{k}}_{i=1}\left(\widebar{X}^{(k)}_{ij}\widebar{X}^{(k)}_{il}-\rho^{(k)}_{jl}\right)\right]\right\}\right]$
$\leq\mathbb{E}\left[\exp\left\{\lambda\tau_{\infty}\left[\sum^{n_{k}}_{i=1}\left(\widebar{X}^{(k)}_{ij}\widebar{X}^{(k)}_{il}-\rho^{(k)}_{jl}\right)\right]\right\}\right]$
@@
when $|\lambda|<2/(\tau_{\infty}\phi)$. Using Proposition 2.9 in Wainwright [2019], for $t\geq 0$, we have that
@@
Consider a zero mean random vector $X=(X_{1},\ldots,X_{d})^{\top}$ with covariance $\Sigma^{\star}$ such that each $X_{j}/\sqrt{\Sigma^{\star}_{jj}}$ is sub-Gaussian with parameter $\sigma$ as defined in Definition B.1. Let $\{X(i)\}^{n}_{i=1}$ be $n$ i.i.d. copies of $X$ and let $\widehat{\Sigma}_{n}=\frac{1}{n}\sum^{n}_{i=1}X(i)X(i)^{\top}$. For $\delta\in(0,1]$, we have
@@
One can view $f$ as a function of $\text{vec}(\Omega)$. Note that $\nabla^{2}f(\text{vec}(\Omega))=\Omega^{-1}\otimes\Omega^{-1}$. The rest of the proof then follows Taylor’s Theorem [Nocedal and Wright, 2006, Theorem 2.1]. ∎
@@
It follows from Example 15.16 in Wainwright [2019] that there exists a ${1}/{2}$-packing of $\mathbb{S}^{p}(s)$, which we denote as $\{\tilde{\theta}^{0},\tilde{\theta}^{1},\ldots,\tilde{\theta}^{M}\}$, such that
@@
which motivates us to compute $\log\det\left(\Omega\right)$ and $\text{tr}\left(\Omega^{-1}\right)$. Let $B=UDV^{\top}$ be the singular value decomposition of $B$ with $U,D,V\in\mathbb{R}^{m\times m}$, $U^{\top}U=UU^{\top}=I_{m}$, $V^{\top}V=VV^{\top}=I_{m}$, and $D=\text{diag}(\lambda_{1},\ldots,\lambda_{m})$. We have the following identities.
@@
Since $\lambda^{2}_{j}\leq\sum^{d}_{i=1}\lambda^{2}_{i}=\|B\|^{2}_{\mathrm{F}}\leq{1}/{16}$, we have $-\lambda^{2}_{j}\geq-{1}/{16}$. Also note that $\log(1+x)\leq x$ for all $x\geq-1$. We thus have
@@
Using Section 9.1.2 of Petersen et al. [2008], we have $\det\left(\Omega\right)=\det\left(I_{m}-B^{\top}B\right)$. Since
@@
Using Section 9.1.3 of Petersen et al. [2008], we have
@@
Appendix C Proof of Theorem 1
We adopt the proof strategy laid out in Wainwright [2019, Chapter 9].
C.1 Additional Notation
@@
[…]
@@
Clearly, for any $\Theta\in\mathbb{M}$ and $\Theta^{\prime}\in\mathbb{M}^{\bot}$, we have that $\langle\Theta,\Theta^{\prime}\rangle=0$
@@ C.3 Remaining proof
In addition, combining Lemma 8 and the fact that $\left[\Theta^{\star}\right]_{\mathbb{M}^{\bot}}=0$ with Wainwright [2019, Lemma 9.14], we have
@@
C.4 Proof of Useful Lemmas
@@ C.4.2 Proof of Lemma 9
[…]
@@ C.4.3 Proof of Lemma 10
=$=\left\{\left\|\Omega^{(k)}+t_{k}\left(\Delta_{\Omega}+\Delta_{\Gamma^{(k)}}\right)\right\|_{2}\right\}^{-2}$
@@ C.4.4 Proof of Lemma 11
Appendix D Theorem 7 and Its Proof
The following theorem provides a high probability error bound for the D-Trace Loss estimator.
Theorem 7.
@@ Appendix D Theorem 7 and Its Proof
The D-Trace Loss estimator is a special case of the FuDGE estimator proposed in Zhao et al. [2019, 2022]. We adapt the proof of Theorem 10 in Zhao et al. [2022] here.
@@
where $\kappa^{(k)}_{\mathcal{L}}$ is the restricted convexity parameter defined in the proof of Theorem 10 of Zhao et al. [2022]. Set
@@
Following the proof of Theorem 10 of Zhao et al. [2022], we have
@@
Appendix E Proof of Theorem 4
Let $R=\lfloor s/d\rfloor$ and let $r=s-Rd$. We follow the construction of the hard-case collection of precision matrices in Step 2 of Proof of Theorem 4.1 in Cai et al. [2016b] to get $\mathcal{F}_{\star}$. We set $M_{n,p}$ and $c_{n,p}$ therein as $M_{n,p}=\text{some constant}$ and $c_{n,p}=R$. By (4.13) and (4.15) in Cai et al. [2016b], for any $\Omega\in\mathcal{F}_{\star}$, we have
@@ Appendix E Proof of Theorem 4
Furthermore, we have $|\Omega|_{0}\leq c_{n,p}d\leq s$. Thus, $\mathcal{F}_{\star}\subseteq\mathcal{G}_{1}$. Following Theorem 6.1 of Cai et al. [2016b], we have
@@
Appendix F Proof of Theorem 5
Without loss of generality, we assume that $d$ and $h$ are even positive numbers. Let $m={d}/{2}$. By Lemma 6, there exist $\{b^{0},b^{1},\ldots,b^{M}\}\subseteq\mathbb{R}^{m^{2}}$ such that
@@ Appendix F Proof of Theorem 5
$\gamma_{\min}\left(\Omega^{j}\right)$ $=\gamma_{\min}\left(\begin{bmatrix}I_{m}&0\\ 0&I_{m}\end{bmatrix}+\begin{bmatrix}0&B^{j}\\ (B^{j})^{\top}&0\end{bmatrix}\right)$
@@
Following Section 2.2 and Corollary 2.6 in Tsybakov [2008], combined with (47)–(49), we have
@@
Appendix G Proof of Theorem 6
@@
[…]
@@
it follows by Theorem 4 that
@@
$\mathcal{G}^{\prime\prime}$ =$=\left\{\Omega\in\mathbb{S}^{d\times d}:\,\Omega\succ 0,\,\Omega=I_{d}+\Delta,\,\Delta_{jj}=0\text{ for all }1\leq j\leq d,\right.$
[…]
Given that for any $\Omega\in\mathcal{G}^{\prime\prime}$, it holds that $\left|\Omega^{-1}\right|_{\infty}\leq\|\Omega^{-1}\|_{2}=\{\gamma_{\min}(\Omega)\}^{-1}\leq M_{\Sigma}$, we can deduce that $\mathcal{G}^{\prime\prime}\subseteq\tilde{\mathcal{G}}^{\prime\prime}$. By Theorem 5, when $M_{\Omega},M_{\Sigma}>1$ and
@@
Appendix H Additional Optimization Details
@@
$-\alpha_{k}\Omega^{-1}_{k}+\rho\Omega_{k}-\rho\tilde{C}^{(t-1)}_{k}=0.$ (53)
The matrix $\Omega^{(t)}_{k}$ is obtained by finding $\Omega_{k}\succ 0$ that satisfies (53). Let $\rho\tilde{C}^{(t-1)}_{k}=U\Lambda U^{\top}$, $\Lambda=\text{diag}\left(\left\{\lambda_{i}\right\}^{d}_{i=1}\right)$, be the eigenvalue decomposition of $\rho\tilde{C}^{(t-1)}_{k}$. Following [Boyd et al., 2011, Section 6.5], we have
@@
$x^{(r)}$ $=\arg\min_{x}\left\{\frac{\rho}{2}\sum^{K}_{k=0}\left(x+y^{(r-1)}_{k}-c_{k}\right)^{2}+\lambda_{\text{M}}|x|+\lambda_{\text{M}}\sum^{K}_{k=0}\sqrt{\alpha_{k}}\left|y^{(r-1)}_{k}\right|\right\}$
@@
Following [Boyd et al., 2011, Section 3.3.1], let $\epsilon^{\text{abs}}>0$ be an absolute tolerance and $\epsilon^{\text{rel}}>0$ be a relative tolerance, we then define the feasibility tolerance for primal feasibility condition $\epsilon^{\text{pri}}>0$ and the feasibility tolerance for dual feasibility condition $\epsilon^{\text{dual}}>0$ at iteration $t$ as
@@
For the proximal gradient descent algorithm to solve D-Trace loss objective (7), the stopping criterion is introduced as below and the detailed algorithm is described in Algorithm 3.
@@
$r^{\text{sub}}$ $=\rho\left|\sum^{K}_{k=0}\left(y^{(r)}_{k}-y^{(r-1)}_{k}\right)\right|,$
$\epsilon^{\text{sub}}$ $=(K+1)\epsilon^{\text{abs}}+\epsilon^{\text{rel}}\rho\max\left\{\sum^{K}_{k=0}\left|y^{(r)}_{k}\right|\,,\,\sum^{K}_{k=0}\left|y^{(r-1)}_{k}\right|\right\}.$
@@
Appendix I Minimax Optimal Rate for Differential Network Estimation
Differential network estimation aims to estimate the difference between two precision matrices without first estimating the individual precision matrices. Although existing studies focus on providing upper bounds for this problem [Zhao et al., 2014, Yuan et al., 2017, Zhao et al., 2022], there is no known matching lower bound, making the minimax optimal rate an open question. As a byproduct of our analysis, in this section, we provide a minimax optimal rate for differential network estimation problem under certain conditions. To the best of our knowledge, this is the first minimax optimal guarantee towards this direction.
@@ Appendix I Minimax Optimal Rate for Differential Network Estimation
where $\mathcal{G}_{2}$ is defined in (23). Now that samples from population $X$ are useless for estimating $\Omega_{Y}$, and we can only rely on samples from population $Y$ to estimate $\Omega_{Y}$. By Theorem 5, we then have
@@
Appendix J Upper Bound for Expected Error
@@
Refer to the proof in Appendix K. The rate described in Theorem 11 consists of two parts. The first part, which is of the order $(s\log d)/{N}$, refers to the estimation of the shared component. The second part, of the order $(h\log d)/{\widebar{n}}$, relates to the estimation of the individual components.
@@
The estimation error in (65) is comprised of three parts: shared component estimation, individual component estimation, and differential network estimation. If $\widebar{n}\geq n_{0}$ and $M_{\Gamma}$ is bounded by a universal constant, the error scales as $\frac{s\log d}{N}+\frac{h\log d}{n_{0}}$. When using only target samples, the lowest error rate achievable is $\frac{(s+h)\log d}{n_{0}}$ as stated in Theorem 4. Therefore, if $N\gg n_{0}$, the error rate can be significantly reduced compared to the optimal rate obtained with only the target samples. Moreover, as demonstrated in Section 5.3, the rate $\frac{s\log d}{N}+\frac{h\log d}{n_{0}}$ is minimax optimal under certain conditions.
Appendix K Proof of Theorem 11
@@
we have $\mathbb{P}\left\{\widebar{\mathbb{G}}(\lambda_{\text{M}})\right\}\leq d^{-\tau^{\prime}}$. Besides, by Section C.3, when (66)–(67) are true, $\mathbb{G}(\lambda_{\text{M}})$ then implies that
@@
=$=\mathbb{P}\left\{\mathbb{G}(\lambda_{\text{M}})\right\}\mathbb{E}\left[\left.\sum^{K}_{k=0}\alpha_{k}\left\|\widecheck{\Omega}^{(k)}-\Omega^{(k)}\right\|^{2}_{\mathrm{F}}\right|\mathbb{G}(\lambda_{\text{M}})\right]+\mathbb{P}\left\{\widebar{\mathbb{G}}(\lambda_{\text{M}})\right\}\mathbb{E}\left[\left.\sum^{K}_{k=0}\alpha_{k}\left\|\widecheck{\Omega}^{(k)}-\Omega^{(k)}\right\|^{2}_{\mathrm{F}}\right|\widebar{\mathbb{G}}(\lambda_{\text{M}})\right].$

### v2

@@ Appendix F Preliminary Lemmas
Definition F.1 (Sub-Gaussian Random Variable, Definition 2 of (8)).
@@
Assume that we obtain samples from $K$ distributions, each with a mean of zero and a covariance matrix of $\Sigma^{(k)}$. Let $\{X^{(k)}_{i}=(X^{(k)}_{i1},\ldots,X^{(k)}_{id})^{\top}\}_{i=1}^{n_{k}}$ represent $n_{k}$ independently distributed samples from the $k$-th distribution. In addition, we assume that $X^{(k)}_{ij}/\sqrt{\Sigma^{(k)}_{jj}}$ is sub-Gaussian with parameter $\sigma$ as defined in Definition F.1. Let
@@
Let $\widebar{X}^{(k)}_{ij}={X^{(k)}_{ij}}/{\sqrt{\Sigma^{(k)}_{jj}}}$, $i=1,\ldots,n_{k}$, $k=1,\ldots,K$, $j=1,\ldots,d$. Define $U^{(k)}_{i,jl}=\widebar{X}^{(k)}_{ij}+\widebar{X}^{(k)}_{il}$, $V^{(k)}_{i,jl}=\widebar{X}^{(k)}_{ij}-\widebar{X}^{(k)}_{il}$, and $\rho^{(k)}_{jl}={\Sigma^{(k)}_{jl}}/{\sqrt{\Sigma^{(k)}_{jj}\cdot\Sigma^{(k)}_{ll}}}$. Then it follows from Proposition 2.9 and (2.18) in Wainwright (10), that
@@
[…]
@@
[…]
@@
when $|\lambda|<2/(\tau_{\infty}\phi)$. Using Proposition 2.9 in Wainwright (10), for $t\geq 0$, we have that
@@
Consider a zero mean random vector $X=(X_{1},\ldots,X_{d})^{\top}$ with covariance $\Sigma^{\star}$ such that each $X_{j}/\sqrt{\Sigma^{\star}_{jj}}$ is sub-Gaussian with parameter $\sigma$ as defined in Definition F.1. Let $\{X(i)\}^{n}_{i=1}$ be $n$ i.i.d. copies of $X$ and let $\widehat{\Sigma}_{n}=\frac{1}{n}\sum^{n}_{i=1}X(i)X(i)^{\top}$. For $\delta\in(0,1]$, we have
@@
One can view $f$ as a function of $\text{vec}(\Omega)$. Note that $\nabla^{2}f(\text{vec}(\Omega))=\Omega^{-1}\otimes\Omega^{-1}$. The rest of the proof then follows Taylor’s Theorem (6, Theorem 2.1). ∎
@@
It follows from Example 15.16 in Wainwright (10) that there exists a 1/2{1}/{2}-packing of $\mathbb{S}^{p}(s)$, which we denote as $\{\tilde{\theta}^{0},\tilde{\theta}^{1},\ldots,\tilde{\theta}^{M}\}$, such that
@@
which motivates us to compute $\log\det\left(\Omega\right)$ and $\text{tr}\left(\Omega^{-1}\right)$. Let $B=UDV^{\top}$ be the singular value decomposition of $B$ with $D,V\in\mathbb{R}^{m\times m}, UU$$^{\top}U=UU^{\top}=I_{m}$, $V^{\top}V=VV^{\top}=I_{m}$, and $D=\text{diag}(\lambda_{1},\ldots,\lambda_{m})$. We have the following identities.
@@
Since $\lambda^{2}_{j}\leq\sum^{d}_{i=1}\lambda^{2}_{i}=\|B\|^{2}_{\mathrm{F}}\leq{1}/{16}$, we have $\lambda^{2}_{j}\geq-{1}/{16}$. Also note that $\log(1+x)\leq x$ for all $x\geq-1$. We thus have
@@
Using Section 9.1.2 of Petersen et al. (7), we have $\det\left(\Omega\right)=\det\left(I_{m}-B^{\top}B\right)$. Since
@@
Using Section 9.1.3 of Petersen et al. (7), we have
@@
Appendix G Proof of Theorem 1
We adopt the proof strategy laid out in Wainwright (10, Chapter 9).
G.1 Additional Notation
@@
The space $\mathbb{H}$ endowed with the inner product $\langle\cdot,\cdot\rangle$ is indeed a Hilbert space. Correspondingly, we have
@@
Clearly, for any $\Theta\in\mathbb{M}$ and $\Theta^{\prime}\in\mathbb{M}^{\bot}$, we have that $\langle\Theta,\Theta^{\prime}\rangle=0$ .
@@ G.3 Remaining proof
In addition, combining Lemma 11 and the fact that $\left[\Theta^{\star}\right]_{\mathbb{M}^{\bot}}=0$ with Wainwright (10, Lemma 9.14), we have
@@
G.4 Proof of Useful Lemmas
@@ G.4.2 Proof of Lemma 12
$\leq|\mathcal{S}_{\Omega}|^{\frac{1}{2}}\left\|\left[\Omega\right]_{\mathcal{S}_{\Omega}}\right\|_{\mathrm{F}}+\sum^{K}_{k=0}\sqrt{\alpha_{k}}\left|\mathcal{S}_{\Gamma^{(k)}}\right|^{\frac{1}{2}}\left\|\left[\Gamma^{(k)}\right]_{\mathcal{S}_{\Gamma^{(k)}}}\right\|_{\mathrm{F}}$ $(\text{Jensen's inequality})$
@@ G.4.3 Proof of Lemma 13
$=\left\{\left\|\Omega^{(k)}+t_{k}\left(\Delta_{\Omega}+\Delta_{\Gamma^{(k)}}\right)\right\|_{2}\right\}^{-2}$
@@ G.4.4 Proof of Lemma 14
Appendix H Corollary 3 and Its Proof
The following corollary of Theorem 10 in Zhao et al. (12) provides a high-probability error bound for the D-Trace Loss estimator.
Corollary 3.
@@ Appendix H Corollary 3 and Its Proof
The D-Trace Loss estimator is a special case of the FuDGE estimator proposed in Zhao et al. (13, 12). We adapt the proof of Theorem 10 in Zhao et al. (12) here.
@@
where $\kappa^{(k)}_{\mathcal{L}}$ is the restricted convexity parameter defined in the proof of Theorem 10 of (12). Set
@@
Following the proof of Theorem 10 of (12), we have
@@
Appendix I Proof of Lemma 2
Let $R=\lfloor s/d\rfloor$ and let r=sRdr=s-Rd. We follow the construction of the hard-case collection of precision matrices in Step 2 of Proof of Theorem 4.1 in Cai et al. (2) to get $\mathcal{F}_{\star}$. We set $M_{n,p}$ and $c_{n,p}$ therein as $M_{n,p}=\text{some constant}$ and $c_{n,p}=R$. By (4.13) and (4.15) in (2), for any $\Omega\in\mathcal{F}_{\star}$, we have
@@ Appendix I Proof of Lemma 2
Furthermore, we have $|\Omega|_{0}\leq c_{n,p}d\leq s$. Thus, $\mathcal{F}_{\star}\subseteq\mathcal{G}_{1}$. Following Theorem 6.1 of (2), we have
@@
Appendix J Proof of Lemma 3
Without loss of generality, we assume that $d$ and $h$ are even positive numbers. Let m=d/2m={d}/{2}. By Lemma 9, there exist $\{b^{0},b^{1},\ldots,b^{M}\}\subseteq\mathbb{R}^{m^{2}}$ such that
@@
[…]
@@
Following Section 2.2 and Corollary 2.6 in Tsybakov (9), combined with (46)–(48), we have
@@
Appendix K Proof of Theorem 2
@@
$\mathcal{G}^{\prime}=\left\{\Omega\in\mathbb{R}^{d\times d}:\Omega\succ 0,\,\left|\Omega\right|_{0}\leq s,\,0<\frac{1}{M_{\Sigma}}\leq\gamma_{\min}\left(\Omega\right)\leq\gamma_{\max}\left(\Omega\right)\leq M_{\Omega}<\infty\right\}.$
@@
it follows by Lemma 2 that
@@
$\mathcal{G}^{\prime\prime}$ $=\left\{\Omega\in\mathbb{S}^{d\times d}:\,\Omega\succ 0,\,\Omega=I_{d}+\Delta,\,\Delta_{jj}=0\text{ for all }1\leq j\leq d,\right$.
$\quad\quad\quad\quad\left.\left|\Delta\right|_{0}\leq h,\,\left|\Delta\right|_{1}\leq M_{\Gamma},\,0<\frac{1}{M_{\Sigma}}\leq\gamma_{\min}(\Omega)\leq\gamma_{\max}(\Omega)\leq M_{\Omega}<\infty\right\}\,.$
Given that for any $\Omega\in\mathcal{G}^{\prime\prime}$, it holds that $\left|\Omega^{-1}\right|_{\infty}\leq\|\Omega^{-1}\|_{2}=\{\gamma_{\min}(\Omega)\}^{-1}\leq M_{\Sigma}$, we can deduce that $\mathcal{G}^{\prime\prime}\subseteq\tilde{\mathcal{G}}^{\prime\prime}$. By Lemma 3, when $M_{\Omega},M_{\Sigma}>1$ and
@@
Appendix L Additional Optimization Details
@@
$\alpha_{k}\Omega^{-1}_{k}+\rho\Omega_{k}-\rho\tilde{C}^{(t-1)}_{k}=0.$ (52)
The matrix $\Omega^{(t)}_{k}$ is obtained by finding $\Omega_{k}\succ 0$ that satisfies (52). Let $\rho\tilde{C}^{(t-1)}_{k}=U\Lambda U^{\top}$, $\Lambda=\text{diag}\left(\left\{\lambda_{i}\right\}^{d}_{i=1}\right)$, be the eigenvalue decomposition of $\rho\tilde{C}^{(t-1)}_{k}$. Following (1, Section 6.5), we have
@@
[…]
@@
Following (1, Section 3.3.1), let $\epsilon^{\text{abs}}>0$ be an absolute tolerance and $\epsilon^{\text{rel}}>0$ be a relative tolerance, we then define the feasibility tolerance for primal feasibility condition $\epsilon^{\text{pri}}>0$ and the feasibility tolerance for dual feasibility condition $\epsilon^{\text{dual}}>0$ at iteration $t$ as
@@
For the proximal gradient descent algorithm to solve D-Trace loss objective (6), the stopping criterion is introduced as below, and the detailed algorithm is described in Algorithm 3.
@@
[…]
@@
Appendix M Minimax Optimal Rate for Differential Network Estimation
Differential network estimation aims to estimate the difference between two precision matrices without first estimating the individual precision matrices. Although existing studies focus on providing upper bounds for this problem (14; 11; 12), there is no known matching lower bound, making the minimax optimal rate an open question. As a byproduct of our analysis, in this section, we provide a minimax optimal rate for differential network estimation problem under certain conditions. To the best of our knowledge, this is the first minimax optimal guarantee towards this direction.
@@
where $\mathcal{G}_{2}$ is defined in (22). Now that samples from population $X$ are useless for estimating $\Omega_{Y}$, and we can only rely on samples from population $Y$ to estimate $\Omega_{Y}$. By Lemma 3, we then have
@@
Appendix N Upper Bound for Expected Error
@@
Refer to the proof in Appendix O. The rate described in Theorem 6 consists of two parts. The first part, which is of the order $(s\log d)/{N}$, refers to the estimation of the shared component. The second part, of the order $(h\log d)/{\widebar{n}}$, relates to the estimation of the individual components.
@@
The estimation error in (64) is comprised of three parts: shared component estimation, individual component estimation, and differential network estimation. If $\widebar{n}\geq n_{0}$ and $M_{\Gamma}$ is bounded by a universal constant, the error scales as $\frac{s\log d}{N}+\frac{h\log d}{n_{0}}$. When using only target samples, the lowest error rate achievable is $\frac{(s+h)\log d}{n_{0}}$ as stated in Lemma 2. Therefore, if $N\gg n_{0}$, the error rate can be significantly reduced compared to the optimal rate obtained with only the target samples. Moreover, as demonstrated in Section 5.3, the rate $\frac{s\log d}{N}+\frac{h\log d}{n_{0}}$ is minimax optimal under certain conditions.
Appendix O Proof of Theorem 6
@@
we have $\mathbb{P}\left\{\widebar{\mathbb{G}}(\lambda_{\text{M}})\right\}\leq d^{-\tau^{\prime}}$. Besides, by Section G.3, when (65)–(66) are true, $\mathbb{G}(\lambda_{\text{M}})$ then implies that
@@
$=\mathbb{P}\left\{\mathbb{G}(\lambda_{\text{M}})\right\}\mathbb{E}\left[\left.\sum^{K}_{k=0}\alpha_{k}\left\|\widecheck{\Omega}^{(k)}-\Omega^{(k)}\right\|^{2}_{\mathrm{F}}\right|\mathbb{G}(\lambda_{\text{M}})\right]+\mathbb{P}\left\{\widebar{\mathbb{G}}(\lambda_{\text{M}})\right\}\mathbb{E}\left[\left.\sum^{K}_{k=0}\alpha_{k}\left\|\widecheck{\Omega}^{(k)}-\Omega^{(k)}\right\|^{2}_{\mathrm{F}}\right|\widebar{\mathbb{G}}(\lambda_{\text{M}})\right].$